### v1

# Data-Efficient Quantum Noise Modeling via Machine Learning

Abstract
Maximizing the computational utility of near-term quantum processors requires predictive noise models that inform robust, noise-aware compilation and error mitigation. Conventional models often fail to capture the complex error dynamics of real hardware or require prohibitive characterization overhead. We introduce a data-efficient, machine learning-based framework to construct accurate, parameterized noise models for superconducting quantum processors. Our approach circumvents costly characterization protocols by learning hardware-specific error parameters directly from the measurement data of existing application and benchmark circuits. The generality and robustness of the framework are demonstrated through comprehensive benchmarking across multiple quantum devices and algorithms. Crucially, we show that a model trained exclusively on small-scale circuits accurately predicts the behavior of larger validation circuits. Our data-efficient approach achieves up to a 65% improvement in model fidelity quantified by the Hellinger distance between predicted and experimental circuit output distributions, compared to standard noise models derived from device properties. This work establishes a practical paradigm for noise characterization, providing a crucial tool for developing more effective noise-aware compilation and error-mitigation strategies.
I Introduction
Quantum computing offers a significant opportunity to address complex optimization problems. In the current noisy intermediate-scale quantum era [1], variational algorithms exemplified by the quantum approximate optimization algorithm (QAOA) [2] and the variational quantum eigensolver (VQE) [3] have been explored for diverse applications, including portfolio optimization [4], unit commitment in power systems [5], and quantum chemistry [6]. Superconducting circuits [7, 8] have emerged as a leading platform because they benefit from established microchip manufacturing techniques and offer fast operating times. However, the practical utility of these devices is constrained by noise mechanisms and hardware limitations that hinder algorithm’s performance. To enhance the reliability of quantum applications on such devices, noise-aware compilation has become an essential strategy. Quantum compilation translates quantum algorithms into sequences of physical operations that can be executed on a target backend. This process ensures that algorithms satisfy the hardware’s intrinsic constraints and optimizes the circuit to minimize the impact of errors. Specifically, the compilation needs to adapt the algorithm to the device’s limited qubit connectivity and small set of supported gates by inserting SWAP gates and decomposing operations in algorithms into the device’s native gate set. This is enabled by compilers supported by, e.g., Qiskit [9], TKET [10] , and Munich quantum toolkit [11, 12]. Beyond ensuring basic executability, noise-aware compilation aims to maximize the fidelity of the compiled circuit. To achieve this, the compiler needs an accurate model of the device’s error characteristics, which often vary over time [13]. This model guides strategic decisions, including selecting a subset of high-fidelity qubits for execution, optimizing the routing of logical qubits to minimize noise accumulation [14, 15], and choosing between different gate sequences that, while logically equivalent, may exhibit varying resilience to noise [16] and other imperfections like Trotter errors [17, 18, 19].
An accurate and predictive noise model is essential for achieving effective noise-aware compilation. However, constructing error models for real hardware presents a significant trade-off between model fidelity and experimental cost. Traditional error models, such as those offered by hardware vendors, typically assume independent Pauli noise channels that fail to capture the intricate dynamics inherent to real hardware. In practice, superconducting quantum devices exhibit spatially and temporally correlated errors [20, 21, 22], non-Markovian memory effects [23, 24, 25], and time-varying decoherence parameters [26]. Conversely, while comprehensive characterization techniques like quantum process tomography [27] can capture these intricate effects, they are resource-intensive and impractical to scale for routine calibration of larger quantum processors. Various intermediate strategies have been explored to bridge this gap. For instance, some approaches use optimization techniques such as genetic algorithms to build models for specific error types [28], while others propose scalable Markovian models that often require custom hardware tuning to be accurate [29]. Further studies have shown that noise models should be tailored to specific applications, as no single model captures all noise dynamics [30]. Even integrated approaches that combine Markovian and non-Markovian effects to improve accuracy still tend to rely on resource-intensive characterizations [31]. These studies highlight the persistent trade-off between model fidelity and experimental overhead, emphasising the necessity for a more efficient approach.
To address these limitations, we introduce a practical and data-efficient alternative to resource-intensive characterization protocols. Our primary contribution is the development of a novel, circuit-size-independent parameterized noise model that achieves high fidelity by leveraging readily available experimental data, such as routine calibration measurements and standard benchmarking results, eliminating the need for dedicated quantum tomography. This approach dramatically reduces the experimental overhead typically associated with building high-fidelity models. As shown in Fig. 1, our data-efficient framework refines an initial parameterized noise model $\mathcal{N}(\bm{\theta})$ through an iterative, machine learning-driven optimization. The objective is to find optimal parameters $\bm{\theta^{*}}$ by minimizing differences between simulated and experimental output distributions, which can be quantified by the Hellinger distance [32]. A crucial feature of this approach is its ability to repurpose training data from any existing circuit execution, such as application benchmarks or algorithm runs. This eliminates the overhead of dedicated characterization protocols and allows the noise model to be tailored specifically to the context of the executed algorithms, ensuring higher fidelity for practical applications. Through extensive benchmarking across diverse quantum circuits and IBM quantum processors, we demonstrate that our optimized noise models significantly outperform the default hardware noise model.
Our framework’s predictive power and scalability are demonstrated through its performance on out-of-distribution validation circuits. Trained solely on small-scale benchmark circuits (4–6 qubits), the model accurately predicts the behavior of larger and more complex circuits (7–9 qubits). This predictive capability delivers a substantial improvement in model fidelity, reducing the mean Hellinger distance to experimental results by 50% and achieving a peak reduction of 65%. This level of predictive accuracy on larger circuits represents a significant advancement over the performance demonstrated in foundational studies on smaller systems [28]. This capability highlights the ability to learn a robust, predictive model from a small, experimentally inexpensive training set, which is critical for any scalable noise characterization protocol. The demonstrated efficacy of our framework establishes a new paradigm for noise characterization where accuracy is decoupled from high experimental overhead. Practitioners can leverage their existing experimental datasets to continuously refine noise models, enhancing the development of noise-aware compilation and error mitigation protocols [33, 34, 35] without incurring additional experimental costs.
Figure 1: Flowchart of the data-efficient quantum noise modeling framework. A parameterized noise model $\mathcal{N}(\bm{\theta})$ is initially constructed based on calibration data provided by hardware vendors. The device-specific noise model $\mathcal{N}(\bm{\theta^{*}})$ is then identified by executing benchmarking circuits on real quantum hardware and aligning experimental results with noisy simulations. This iterative optimization, performed on a classical computer, leverages machine learning (ML) to update $\bm{\theta}$. This approach is highly flexible, designed to ingest outcome distributions from diverse preexisting experimental data, including application-specific circuits, avoiding the need for dedicated characterization.
The paper is structured as follows. Section II presents the theoretical framework and methodology, establishing the baseline provided by device-calibrated noise models and introducing our parameterized noise model. It also details the ML technique employed for parameter optimization and concludes with a critical analysis of the framework’s scalability. Section III demonstrates a thorough experimental validation of our approach, benchmarking the model’s predictive accuracy against data from multiple superconducting processors using a diverse set of applications, including the QAOA [2], VQE [3], and random circuits. This validation includes an in-depth single-device characterization, a cross-platform comparative analysis to demonstrate hardware generality, and a study of training efficiency. Finally, Sec. IV concludes the paper.
II Theoretical Framework and Methodology
This section details our noise modeling methodology. We first introduce a baseline noise model for comparison. We then introduce our circuit-size-independent parameterized noise model, followed by optimization strategy. Finally, we discuss the scalability of our approach.
II.1 Default noise model
Most providers of superconducting quantum processing units (QPUs) offer noise models, which are often derived directly from the quantum device’s calibration data. A common example is the standard model available in Qiskit [9] for IBM devices, which can be constructed using the provided backend configurations and properties. This model isolates errors into two primary categories: gate errors and readout errors. Gate errors are modeled as a composed quantum channel $\mathcal{E}_{\text{gate}}$ applied after each ideal unitary operation. This channel is formed by the composition of a depolarizing channel $\mathcal{E}_{\text{depol}}$, followed by a thermal relaxation channel $\mathcal{E}_{\text{thermal}}$:
$\mathcal{E}_{\text{gate}}=\mathcal{E}_{\text{thermal}}\circ\mathcal{E}_{\text{depol}},$ (1)
where the composition $\mathcal{A}\circ\mathcal{B}$ implies that channel $\mathcal{B}$ is applied first. The thermal relaxation channel $\mathcal{E}_{\text{thermal}}$ models energy relaxation ($T_{1}$) and dephasing ($T_{2}$) processes occurring during the gate’s finite duration. The depolarizing channel $\mathcal{E}_{\text{depol}}$ is not an independently measured quantity; rather, its error probability is chosen such that the infidelity of the composite channel $\mathcal{E}_{\text{gate}}$ precisely matches the experimentally characterized gate error rate $e_{g}$.
[…]
Although this baseline model can be easily obtained, it excludes cross-talk and other correlated effects. Moreover, it captures only noise mechanisms explicitly characterized during routine device calibration procedures, creating a static model that cannot account for the subsequent temporal evolution of the physical device. To better capture the real noise behavior of quantum algorithms executing on QPUs, we develop a tractable, data-efficient error model, which we detail in the following section.
II.2 Parameterized noise model
The model is designed to be parameterized by a vector of learnable coefficients $\bm{\theta}$, which are chosen using ML techniques to minimize the discrepancy between simulation and experimental data. An error channel $\mathcal{E}_{g}$ for any gate operation is modeled as the composition of several distinct error channels, each targeting a specific physical error mechanism.
II.2.1 Single-qubit gate errors
The error channel $\mathcal{E}_{\text{1Q}}$ for any single-qubit gate (e.g., $X$, SXSX) is modeled as a composition of four channels:
$\mathcal{E}_{\text{1Q}}=\mathcal{E}_{\text{ph}}\circ\mathcal{E}_{\text{amp}}\circ\mathcal{E}_{\text{dep}}\circ\mathcal{E}_{\text{coh}}.$ (2)
The coherent error channel $\mathcal{E}_{\text{coh}}$ models systematic over- or under-rotations due to control pulse distortions. It is represented by a unitary operator $U_{\text{coh}}=\exp(-iH_{\text{coh}}t_{g})$, where $t_{g}$ is gate duration and the error Hamiltonian $H_{\text{coh}}$ is a linear combination of Pauli operators with learnable amplitudes $\theta_{x,y,z}$:
$H_{\mathrm{coh}}=\theta_{x}X+\theta_{y}Y+\theta_{z}Z.$ (3)
The channel acts as $\mathcal{E}_{\text{coh}}(\rho)=U_{\text{coh}}~\rho~U_{\text{coh}}^{\dagger}$. For the identity gate, this coherent error is omitted.
The depolarizing error channel $\mathcal{E}_{\text{dep}}$ models a generic, incoherent disruption of the quantum state. The single-qubit depolarizing channel is given by
$\mathcal{E}_{\text{dep}}(\rho)=(1-\lambda_{\text{dep}})\rho+\frac{\lambda_{\text{dep}}}{3}(X\rho X+Y\rho Y+Z\rho Z).$ (4)
The depolarization probability $\lambda_{\text{dep}}$ is derived from the hardware-reported gate error $e_{g}$ and gate duration $t_{g}$ via a duration-aware, saturating function
$\lambda_{\text{dep}}(e_{g},t_{g})=1-e^{-(k_{\text{dep}}e_{g}+b_{\text{dep}})t_{g}/t_{\text{char}}}.$ (5)
Here, $k_{\text{dep}}$ and $b_{\text{dep}}$ are learnable parameters that scale and offset the error rate, and $t_{\text{char}}$ is a constant characteristic gate time for the device. This form captures the physical reality that error probability saturates for long-duration operations.
The amplitude damping channel $\mathcal{E}_{\text{amp}}$ models energy relaxation. Its probability $\lambda_{\text{amp}}$ is based on the physical relaxation process for a gate of duration $t_{g}$, with an added learnable offset $b_{\text{amp}}$
$\lambda_{\text{amp}}(t_{g},T_{1})=1-e^{-t_{g}/T_{1}}+b_{\text{amp}}.$ (6)
To capture non-Markovian dephasing, potentially arising from low-frequency noise sources like two-level-system fluctuators, we model phase damping channel $\mathcal{E}_{\text{ph}}$ using a stretched exponential decay. The probability of a phase-flip error, $\lambda_{\phi}$, is given by
$\lambda_{\phi}(t_{g},T_{\phi})=1-e^{-2\left(t_{g}/T_{\phi}\right)^{\beta_{\text{1Q}}}},$ (7)
where $T_{\phi}$ is the pure dephasing time derived from the measured $T_{1}$ and $T_{2}$, satisfying $1/T_{2}=1/(2T_{1})+1/T_{\phi}$, and $\beta_{\text{1Q}}$ is a learnable exponent.
II.2.2 Two-qubit gate errors
The error channel for a two-qubit entangling gate, $\mathcal{E}_{\text{2Q}}$, follows a similar compositional structure but includes terms for crosstalk and correlated errors:
$\mathcal{E}_{\text{2Q}}=\mathcal{E}_{{zz}}\circ(\mathcal{E}_{\text{ph}}\otimes\mathcal{E}_{\text{ph}})\circ(\mathcal{E}_{\text{amp}}\otimes\mathcal{E}_{\text{amp}})\circ\mathcal{E}_{\text{dep}}^{(2)}\circ\mathcal{E}_{\text{coh}}^{(2)}.$ (8)
Two-qubit coherent error $\mathcal{E}_{\text{coh}}^{(2)}$ models coherent crosstalk, such as residual $ZZ$ coupling or control crosstalk. The error Hamiltonian includes two-qubit Pauli terms
$H_{\text{coh}}^{(2)}=\theta_{ix}I\otimes X+\theta_{zx}Z\otimes X+\theta_{zz}Z\otimes Z,$ (9)
where $\theta_{ix,zx,zz}$ are learnable parameters. Two-qubit depolarizing channel $\mathcal{E}_{\text{dep}}^{(2)}$ introduces two learnable parameters $k_{\text{dep,2Q}}$ and $b_{\text{dep,2Q}}$, with a constant characteristic gate time $t_{\text{char,2Q}}$, forming the duration-aware, saturating probability $\lambda_{\text{dep}}^{(2)}$ analogous to the single-qubit case in Eq. (5). The single-qubit damping and dephasing channels ($\mathcal{E}_{\text{amp}},\mathcal{E}_{\text{ph}}$) are applied independently to each qubit involved in the gate, using parameters specific to the two-qubit context, including $b_{\text{amp,2Q}}$ $b_{\text{amp,2Q}}$, $\beta_{\text{2Q}}$, and $b_{\phi\text{,2Q}}$. Note that the offset parameter $b_{\phi\text{,2Q}}$ is additionally introduced for two-qubit gates, compared to single qubit gates Eq. (7), to better characterize the more complex error mechanisms of two-qubit gates.
The correlated $ZZ$ dephasing channel $\mathcal{E}_{{zz}}$ introduces a correlated dephasing error, a common mechanism in superconducting qubits arising from spectator qubit interactions or flux noise. It is modeled as a Pauli channel that applies a $Z\otimes Z$ operator with probability $\lambda_{zz}$
$\mathcal{E}_{{zz}}(\rho)=(1-\lambda_{zz})\rho+\lambda_{zz}(Z\otimes Z)\rho(Z\otimes Z).$ (10)
The parameter $\lambda_{zz}$ is also modeled as a saturating, duration-dependent function parameterized by a learnable rate $k_{zz}$
$\lambda_{zz}(t_{g})=1-e^{-k_{zz}t_{g}/t_{\text{char,2Q}}}.$ (11)
II.2.3 Readout errors
[…]
$M_{0,0}\rightarrow M_{0,0}-\lambda_{\text{corr}}\quad\text{and}\quad M_{0,3}\rightarrow M_{0,3}+\lambda_{\text{corr}}.$ (12)
[…]
$\lambda_{\text{corr}}=a~\bar{e}+b,$ (13)
[…]
Table 1: The 20 learnable parameters $\bm{\theta}$ of the noise model, grouped by function role.
Parameter Description Type
Single-Qubit Gate Parameters (7)
$k_{\text{dep}}$ Depolarizing scaling factor Stochastic
$b_{\text{dep}}$ Depolarizing offset Stochastic
$b_{\text{amp}}$ Amplitude damping offset Stochastic
$\theta_{x,y,z}$ Coherent rotation angles Coherent
$\beta_{\text{1Q}}$ Stretched dephasing exponent Non-Markovian
Two-Qubit Gate Parameters (9)
$k_{\text{dep,2Q}}$ Depolarizing scaling factor Stochastic
$b_{\text{dep,2Q}}$ Depolarizing offset Stochastic
$b_{\text{amp,2Q}}$ Amplitude damping offset Stochastic
$b_{\phi\text{,2Q}}$ Phase damping offset Stochastic
$\theta_{ix,zx,zz}$ Coherent crosstalk angles Coherent
$\beta_{\text{2Q}}$ Stretched dephasing exponent Non-Markovian
$k_{zz}$ Correlated $ZZ$ dephasing rate Correlated
Correlated Readout Parameters (4)
$a,b_{00\leftrightarrow 11}$ Coefficients ($|00\rangle\leftrightarrow|11\rangle$) Correlated
$a,b_{01\leftrightarrow 10}$ Coefficients ($|01\rangle\leftrightarrow|10\rangle$) Correlated
The complete noise model, encompassing gate and readout error channels detailed above, is fully specified by a set of 20 free parameters $\bm{\theta}$. These parameters quantify the strength of each physical error mechanism, from coherent rotations and stochastic decoherence to non-Markovian dynamics and correlated readout, forming the target vector for our data-efficient optimization. For clarity, Table 1 summarizes these learnable model parameters, grouped by their functional role: 7 for single-qubit, 9 for two-qubit, and 4 for readout error characterization.
II.3 Parameter optimization
The error model introduced in the previous section contains 20 free parameters $\bm{\theta}$. Our goal is to choose the parameters of the extended model such that the model accurately reproduces sample distributions $P_{i}$ obtained from executing circuits $C_{i}\in\mathcal{C}$ on a specific quantum computing backend. This can be formalized as identifying the parameter configuration $\bm{\theta}^{*}$ that minimizes the Hellinger distance [32] between simulation results and the results obtained on a real device. The Hellinger distance $D_{H}$ ranges from 0 to 1, with 0 indicating identical distributions, and is defined as
$D_{H}(P,Q)=\left(1-\sum_{j}\sqrt{p_{j}q_{j}}\right)^{1/2},$ (14)
where $p_{j}$ and $q_{j}$ are probabilities of outcome $j$ in distributions $P$ and $Q$, respectively. Specifically, we minimize the mean Hellinger distance across all circuits in the training set, with respect to the circuits of interest, to obtain
$\bm{\theta}^{*}=\text{argmin}_{\bm{\theta}}\left\{\frac{1}{|{\mathcal{C}}|}\sum_{i}H[P_{i},Q_{i}(\bm{\theta})]\right\}\,.$ (15)
Here, $|{\mathcal{C}}|$ is the size of the training set, $Q_{i}(\bm{\theta})$ is the empirical distribution obtained by simulating the circuit with a noise model with parameters $\bm{\theta}$. Since the configuration space of the model is large and evaluating the Hellinger distance requires running noisy simulations, identifying $\bm{\theta}^{*}$ is not tractable using brute force search. We thus employ Bayesian optimization (BO) to efficiently obtain estimates of $\bm{\theta}^{*}$. Bayesian optimization is a global optimization technique particularly effective for expensive-to-evaluate black-box objective functions [36]. It employs a probabilistic surrogate model to intelligently select parameter evaluations by maximizing an acquisition function that balances exploration of uncertain regions with exploitation of promising areas.
As a surrogate model, we use a tree-structured Parzen estimator [37], which uses adaptive, non-parametric Parzen window density models to distinguish promising from poor regions of the search space based on past evaluations. Then samples new configurations from the most promising areas to enable efficient hyperparameter optimization in complex, hierarchical domains. We benchmark the BO against random search (RS) in which parameter points are randomly selected from the search space to provide unbiased model estimates. Both methods are implemented using Optuna [38].
II.4 Scalability analysis
A key advantage of our approach is that we determine a fixed set of parameters within the noise model, irrespective of the circuit size or qubit count, leading to a circuit-size-independent complexity. This design choice decouples the computational complexity from the scale of the quantum hardware or the dimensions of the quantum system under study, enabling applicability to large-scale systems without prohibitive resource demands.
The resource requirements of our framework consist of experimental and computational costs. Experimentally, our data-efficient approach circumvents the prohibitive data overhead associated with full process tomography, which scales exponentially with qubit count due to the necessity of exhaustive state or process reconstructions. Instead, it relies solely on circuit-level outcome distributions obtained from a compact, predefined set of benchmark circuits executed on the target quantum hardware. Computationally, the primary cost arises from the classical simulation of these benchmark circuits, required for each evaluation of the objective function during parameter determination. Notably, the cost per determination step remains constant, as it involves simulating the entire fixed set of training circuits to compute the discrepancy against the experimental dataset. Each iteration evaluates an objective function that quantifies this mismatch, here, via the Hellinger distance. The selection of the ML is pivotal for minimizing the number of such evaluations, given their computational intensity. We adopt BO, which excels in handling expensive black-box functions, accelerating convergence and enhancing the framework’s scalability. Moreover, the entire characterization process is conducted as an offline, classical post-processing step, obviating any need for real-time interaction with the QPU.
III Experimental Results
We benchmark our method using data collected from multiple IBM QPUs by executing quantum algorithms with varying qubit counts. The chosen systems ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen comprise 27 qubits and use CNOT gate as their native two-qubit entangling operation, whereas ibm_kingston features 156 qubits and uses CZ gate. The benchmarks include the QAOA and VQE, which are critical classes of variational algorithms designed for near-term hardware. We further include random circuits to evaluate the framework’s suitability for unstructured quantum computations. Moreover, we partition our benchmark circuits by size into a small-size training set (4–6 qubits) and a large-size validation set (7–9 qubits) for testing the model’s ability to extrapolate from smaller systems to predict the behavior of larger, more complex systems.
III.1 QAOA
The QAOA constructs a parameterized quantum circuit consisting of alternating unitaries generated by a problem-dependent cost Hamiltonian $H_{C}$ and a mixing Hamiltonian $H_{M}$:
$U(\bm{\gamma},\bm{\beta})=\prod_{l=1}^{L}e^{-i\beta_{l}H_{M}}e^{-i\gamma_{l}H_{C}},$ (16)
where $\bm{\gamma}$ and $\bm{\beta}$ are variational parameters determined classically to minimize the expectation value of $H_{C}$. The mixing Hamiltonian is commonly chosen as $H_{M}=\sum_{i}X_{i}$, which promotes exploration of the Hilbert space. The cost Hamiltonian encodes the objective function of the optimization problem, typically taking the form of a diagonal operator in the computational basis
$H_{C}=\sum_{i}c_{i}Z_{i}+\sum_{i<j}c_{ij}Z_{i}Z_{j},$ (17)
where $Z_{i}$ are Pauli-$Z$ operators and the coefficients $c_{i},c_{ij}$ reflect the problem structure. In this work, we employ QAOA applied to portfolio optimization problems, which involves fully connected two-qubit gates; further details can be found in [15, 16, 39, 4].
For each QAOA circuit, we use the constrained optimization by linear approximation (COBYLA) [40] to choose angles $\bm{\gamma}$ and $\bm{\beta}$. Additionally, we use SWAP networks [15, 41] on linear topologies to address the connectivity constraints of the QPU, which enables all compiled circuits to share the same fundamental gate structure and qubit connectivity pattern, regardless of the target device. This standardization allows us to control for variability in circuit design and compilation strategies.
We quantify circuit depth and gate counts of benchmark circuits executed on ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen to assess their complexity and noise sensitivity. As detailed in Table 2, the training set comprises circuits with 4–6 qubits and depths of $27.0\pm 3.4$, while the prediction set exhibits 7–9 qubits and depths of $39.0\pm 3.4$. The scaling of CNOT gates, from $27.0\pm 10.7$ in training to $78.0\pm 18.4$ in prediction, reflects the increased entanglement complexity and noise accumulation in larger circuits.
Table 2: Properties of QAOA circuits executed on ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen. Metrics include circuit depth and gate counts (CNOT and single-qubit).
Property Training set Prediction set
Depth $27.0\pm 3.4$ $39.0\pm 3.4$
CNOT gates $27.0\pm 10.7$ $78.0\pm 18.4$
Single-qubit gates $50.3\pm 10.7$ $92.3\pm 13.2$
III.1.1 Single-device characterization
Refer to caption(b)
Figure 2: Hellinger distances between outcomes from simulations using default device noise model ($x$-axis) and our parameterized noise models ($y$-axis) versus real executions on ibmq_kolkata. Blue circles represent models optimized via BO, while orange squares denote those optimized through RS. All points below the dashed diagonal ($y=x$) indicate an improvement in model fidelity over the default model. The percentages shown in the legend represent the average reduction in Hellinger distance achieved by each optimization method relative to the default model. (a) Training phase using QAOA circuits on 4-, 5-, and 6-qubit instances, and (b) predictive performance on 7-, 8-, and 9-qubit circuits. Each circuit size was executed at four distinct times, yielding a total of 12 circuits for both training and prediction sets.
(a)(b)
Figure 3: Optimization landscape of the mean Hellinger distance, projected onto the two most significant noise parameters, as identified by Optuna, with marginalization over other noise model parameters. (a) Bayesian optimization and (b) random search. Each point represents a trial, and the red star marks the optimal configuration yielding the minimum value.
Refer to caption(b)
Figure 4: Comparison of BO and RS for training the parameterized noise model on ibmq_kolkata (Kolkata-RS and Kolkata), alongside BO results on ibmq_mumbai (Mumbai) and ibmq_ehningen (Ehningen). (a) The average percentage reduction in Hellinger distance $D_{H}$ relative to the default noise model, for both training (blue, left) and prediction (orange, right) sets, evaluated on QAOA circuits with 4–6 qubits (training) and 7–9 qubits (prediction), each executed at four calibration instances. Error bars denote means and standard deviations derived from three independent repetitions across the 12 circuits per set. (b) The convergence of the optimization process corresponding to the training results presented in (a), which displays the best-achieved mean Hellinger distance of the training circuits over iterations.
We evaluate the method on ibmq_kolkata. On the training set, our parameterized noise model optimized with BO and RS achieves an average Hellinger distance of 0.095 and 0.118, respectively. This represents a significant reduction in predictive error of 45.92% and 32.50% compared to the default model’s distance of 0.175, as shown in Fig. 2(a). Crucially, the model’s performance generalizes robustly to the validation set (7–9 qubits) (Fig. 2(b)). For these larger circuits, our model achieves an average Hellinger distance of 0.135 (BO) and 0.148 (RS), corresponding to a reduction of 50.91% and 46.22% over the default model’s distance of 0.275. These results also show that BO consistently outperforms RS in both the training and prediction processes, demonstrating its high optimization efficiency.
The contour plots in Fig. 3(a) and 3(b) compare the optimization landscapes explored by BO and RS, respectively, within a two-dimensional parameter space, which represent the most important parameters determined by Optuna. We observe that BO exhibits a structured and adaptive pattern, with sampled points increasingly concentrated in regions of minimal Hellinger distance, demonstrating that it efficiently identifies promising parameter regions. In contrast, RS yields a uniform and inefficient distribution of trials across the parameter space, resulting in a diffuse exploration that fails to converge effectively on the global minimum. This demonstrates the advantage of employing BO to achieve data-efficient parameterization.
III.1.2 Cross-platform comparative analysis
To assess the robustness of our data-efficient framework, we repeat the full training and prediction pipeline on ibmq_kolkata three times, comparing BO to RS and plotting results with error bars in Fig. 4. As shown in Fig. 4(a), BO (Kolkata) outperforms RS (Kolkata-RS) in both training and prediction, exhibiting better performance. Moreover, Fig. 4(b) demonstrates that BO achieves a lower Hellinger distance than RS, confirming faster convergence to high-fidelity noise models.
In the following, we employ BO for determining the parameters of noise models and extend our analysis across two other QPUs (ibmq_mumbai and ibmq_ehningen) featuring varying noise characteristics to evaluate the generalizability and robustness of our approach. For each device, we apply the same parameterized model architecture but train it independently on device-specific data. This process yields a unique model tailored to the specific error profile of each hardware instance. As shown in Fig. 4, our approach consistently yields a substantial reduction in $D_{H}$ across all backends, and the optimization process converges to a low final $D_{H}$ value. The framework’s effectiveness is most pronounced on ibmq_mumbai, which exhibits an average $D_{H}$ reduction of over 50% for both training and prediction sets, with the reduction for individual prediction circuits peaking at 65%. Strong performance is also observed on ibmq_kolkata (45% training, 49% prediction) and ibmq_ehningen (35% training, 32% prediction). We summarize these achievements in Table 3. This high-fidelity performance across multiple systems, each with distinct error characteristics, confirms the robustness and broad applicability of our methodology for capturing hardware-specific noise profiles.
Table 3: Experiment details of QAOA circuits executed on ibmq_kolkata, ibmq_mumbai, and ibmq_ehningen. The circuit fidelity $F$ is defined in Eq. (18). The physical qubits represent the active qubits selected for circuit execution within each benchmark set.
ibmq_kolkata ibmq_mumbai ibmq_ehningen
Property Training Prediction Training Prediction Training Prediction
Default $D_{H}$ $0.175\pm 0.000$ $0.275\pm 0.000$ $0.179\pm 0.001$ $0.277\pm 0.001$ $0.230\pm 0.000$ $0.328\pm 0.000$
Parameterized $D_{H}$ $0.096\pm 0.002$ $141\pm 0.004 0.0$$089\pm 0.001$ $0.137\pm 0.001$ $0.151\pm 0.001$ $0.225\pm 0.004$
$D_{H}$ reduction (%) $45.3\pm 0.8$ $48.5\pm 1.7$ $4\pm 0.7 50.5±0.$$5\pm 0.5$ $34.4\pm 0.6$ $31.3\pm 1.2$
Circuit fidelity $F$ (%) $77.6\pm 8.1$ $47.9\pm 8.3$ $77.4\pm 7.1$ $51.5\pm 6.7$ $79.7\pm 6.9$ $53.2\pm 8.0$
Physical qubits
{12, 15, 18,
21, 23, 24}
{4, 7, 10, 12, 15,
18, 21, 23, 24}
{14, 16, 19, 20,
22, 24, 25}
{12–16, 19,
22, 24, 25}
{15, 18–26} {12–15, 18–26}
Physical qubit count 6 9 7 9 10 13
We use device calibration data to calculate the circuit fidelity $F$ to analyze the noise level of executing circuits on each device:
$F=\prod_{i}(1-\epsilon_{i}),$ (18)
where $\epsilon_{i}$ are gate and readout errors. The increase in complexity from training set to prediction set corresponds to a significant circuit fidelity degradation from about $78\%$ to about $50\%$, as shown in Table 3, underscoring the challenge of extrapolating noise models to larger circuit regimes where errors become more pronounced. Notably, the physical qubit allocation, driven by device-specific transpilation strategies, reveals that prediction circuits span a broader device region compared to training circuits, potentially introducing greater variability from spatial heterogeneity in device error rates. We observe that ibmq_ehningen, compared to other QPUs, maps circuits across a significantly wider physical qubit topology across four different benchmark executions. The increased spatial complexity is probably responsible for the more modest performance gains. A potential reason is that our amount of training data is insufficient to fully characterize the more complex error landscape of an expanded qubit ensemble. Nevertheless, the systematic increase in circuit complexity from training to prediction provides a rigorous test of our noise model’s ability to extrapolate beyond its training regime. The key aspect of this extrapolation is the model’s capacity to characterize qubits absent from the training set. This is enabled because our framework avoids qubit-specific error rate mappings. Instead, it constructs a generative model of the device’s noise, optimizing global hyperparameters $\bm{\theta}$ that parameterize statistical distributions for gate errors from which individual qubit and gate parameters are drawn.
Figure 5: Impact of different training dataset sizes on the performance of BO-optimized noise model on ibmq_kolkata. Each training set is denoted by the number of QAOA circuits. The experimental setting is the same as in Fig. 4.
III.1.3 Training efficiency study
This section examines the interplay between model performance and training dataset size to identify the minimal data volume needed for target accuracy with reduced resource demands. To assess our framework’s data requirements, we train and validate our noise model using four distinct QAOA datasets of varying sizes (9, 12, 15, and 18 circuits). The results reveal a remarkable degree of robustness and data efficiency. Across 9 to 18 distinct QAOA circuit sizes, circuit properties remain remarkably stable. The preserved complexity distribution across all dataset sizes, combined with consistent fidelity estimates, validates that our smallest 9-circuit training set adequately captures the noise characteristics necessary for model optimization; more details can be found in Table 6 in Appendix A. This structural invariance explains the robust performance shown in Fig. 5, where the framework achieves a consistently high reduction in $D_{H}$ of approximately 45% to 50% across all tested datasets, despite a $2\times$ difference in training data volume. Moreover, the optimization convergence profiles for 12, 15, and 18 circuit datasets are nearly indistinguishable, reaching the same optimal $D_{H}$ value at a similar rate, highlighting stability (see Fig. 6 in Appendix A). This key finding confirms that our method does not require extensive, experimentally costly training data to construct a high-fidelity noise model, highlighting its practical efficiency.
III.2 VQE
To verify the efficacy of the method beyond QAOA, we perform a similar optimization for VQE circuits. We consider the task of minimizing the ground state energy of the cluster Hamiltonian
$H_{n}=\alpha_{0}Z_{0}+\alpha_{n}Z_{n}+\sum_{j=0}^{n-3}\beta_{j}Z_{j}X_{j+1}Z_{j+2}\,$, (19)
where $\alpha_{i}$ and $\beta_{i}$ are real coefficients. In the following, we choose $\alpha_{k}=1/n$ ($\beta_{k}=1/n$) for all $k$ for $n=4,\dots,9$. Using a simple, hardware-efficient ansatz corresponding to the EfficientSU2 circuit in Qiskit, we optimize the parameters using the sequential least squares quadratic programming (SLSQP) [42] in a noiseless simulation and verify that the optimized circuits converge to the true ground state energy. We then execute the optimized circuits on the 156-qubit Heron R2 QPU ibm_kingston. For the execution, we group the circuits for each $n$ into two qubit-wise commuting groups and sample from each group with $30.000$ shots. This results in a total of $12$ circuits which we partition into six circuits ($n=4,5,6$) for training and six circuits for validation ($n=7,8,9$).
Performing three independent optimizations using $200$ iterations each, we obtain optimized parameters which result in an improvement of $43.99\%\pm 0.14\%$ in model fidelity for the training set and $18.58\%\pm 0.90\%$ for the validation set (up to 27% for a circuit prediction, see Fig. 7(a) in Appendix A). While the parameterized model aligns more closely with experiment, especially on the training circuits, the improvement on the validation set is less pronounced than in the QAOA case. This likely reflects differences between the devices from which the data were sampled. As detailed in Table 4, VQE circuits contain significantly fewer entangling gates (CZs) than QAOA circuits (CNOTs). Moreover, the ibm_kingston processor itself exhibits lower error rates. These factors contribute to inherently higher baseline circuit fidelities, resulting in substantially smaller Hellinger distances between the baseline model and experiment. On the validation set, the default model’s Hellinger distance is roughly half of that observed for the comparable model on, e.g., ibmq_kolkata. Consequently, the margins of improvement over the baseline are smaller. Nevertheless, these results show that the approach in this work is robust across device generations, qubit counts, and native gate sets, establishing a clear path forward for modeling noise in more complex VQE circuits.
Table 4: Properties of VQE circuits executed on ibm_kingston.
Property Training set Prediction set
Default $D_{H}$ $0.075\pm 0.000$ $0.124\pm 0.000$
Parameterized $D_{H}$ $0.042\pm 0.000$ $0.101\pm 0.001$
$D_{H}$ reduction (%) $43.99\%\pm 0.14\%$ $18.58\%\pm 0.90\%$
Depth $41.2\pm 3.2$ $53.7\pm 3.4$
CZ gates $12.0\pm 2.7$ $21.0\pm 2.7$
Single-qubit gates $107.2\pm 21.9$ $184.7\pm 19.3$
Circuit fidelity $F$ (%) $91.0\pm 1.5$ $83.8\pm 3.9$
Physical qubits
{0–5}
{0–8}
Physical qubit count 6 9
III.3 Random circuits
We further benchmark the method through random circuits on ibmq_kolkata. Random quantum circuits were generated using Qiskit’s random_circuit function with systematically varied parameters: qubit count $n\in\{4,...,10\}$ and depth $d=n+3$. The function constructs circuits by randomly sampling single-qubit gates from SU(2) and placing two-qubit gates between random qubit pairs within the depth constraint, with terminal measurements on all qubits. The results show a substantial mean improvement of 43.60% in model fidelity for training and 45.90% for prediction (see Fig. 7(b) in Appendix A). We perform three independent optimization processes and achieve a mean Hellinger distance reduction of $42.9\%\pm 1.5\%$ for the training set and $46.8\%\pm 2.1\%$ for the prediction set, with a maximum reduction of 63% observed for a circuit in the prediction set. This demonstrates that our method’s high performance is a statistically significant result, making it a reliable tool for practical hardware characterization.
Table 5: Properties of random circuits executed on ibmq_kolkata.
Property Training set Prediction set
Default $D_{H}$ $0.162\pm 0.000$ $0.250\pm 0.000$
Parameterized $D_{H}$ $0.093\pm 0.002$ $0.133\pm 0.005$
$D_{H}$ reduction (%) $42.9\pm 1.3$ $46.8\pm 1.9$
Depth $87.2\pm 18.9$ $182.6\pm 23.8$
CNOT gates $38.3\pm 17.8$ $136.6\pm 20.4$
Single-qubit gates $106.8\pm 32.0$ $271.7\pm 34.3$
Circuit fidelity $F$ (%) $64.0\pm 14.0$ $24.8\pm 7.8$
Physical qubits
{0–7, 10, 12, 15,
17, 18, 21, 23, 24}
{0, 1, 4, 6, 7, 10,
12–19, 21, 23, 24}
Physical qubit count 16 17
[…]
IV Discussion and Conclusion
This work develops a novel parameterized noise model with circuit-size-independent complexity. Leveraging machine learning, we repurpose data from a few routine circuit executions to construct high-fidelity, hardware-specific noise models for superconducting quantum processors. By systematically searching complex, multi-parameter error spaces, our approach consistently produces models that are significantly more predictive than default ones obtained from hardware vendors. We demonstrate a reduction of up to 65% in the Hellinger distance, a metric that quantifies the divergence between predicted and experimental quantum state distributions, confirming that our ML-optimized models produce output distributions that more accurately represent the behavior of algorithms executed on real quantum hardware. This significant improvement in model fidelity is essential for the reliable simulation of quantum circuits on noisy quantum devices, enabling a more accurate noise-aware quantum circuit compilation by providing a precise foundation for optimizing qubit selection, SWAP gate insertion, and other compilation passes.
A key advantage of our approach is its efficiency and practicality. It repurposes existing benchmarking or application-specific execution data, avoiding the need for extensive characterization protocols such as quantum process tomography. This reusability reduces experimental overhead, enabling practitioners to leverage prior executions to improve algorithm testing and error mitigation protocols. Moreover, this allows for the continuous refinement of noise models, ensuring scalability for larger quantum circuits and processors. Comprehensive benchmarking across multiple IBM quantum devices and quantum algorithms validates the framework’s adaptability to different types of superconducting hardware, highlighting its generality and potential for broad adoption in applications. By bridging the gap between theoretical models and experimental reality, our data-efficient approach provides a critical enabling technology for developing the next generation of effective error mitigation techniques and noise-aware compilers, accelerating the realization of practical quantum computing applications.
Future work should balance model complexity and computational cost for scaling to larger quantum systems. More parameters enhance accuracy but increase ML optimization time and data requirements, potentially requiring dedicated characterization circuits to prevent overfitting. A promising route is hierarchical noise modeling that partitions the coupling graph into regional subgraphs with region-specific parameters, capturing spatial inhomogeneity without percomponent overparameterization and preserving scalability. Additionally, extending our framework to model leakage, where qubits excite outside of the computational subspace, will be critical. This requires higher-dimensional simulations and often specialized characterization protocols, substantially increasing the computational burden of the ML loop. Although beyond the scope of this initial study, incorporating leakage is essential for further fidelity gains. The ultimate goal is a full integration into quantum compiler toolchains, enabling adaptive, algorithm-specific compilation that dynamically tailors noise models to both hardware’s current state and circuit structure. This would optimize execution by capturing the unique interplay between algorithms and device-specific error channels. Finally, extending to trapped-ion and photonic platforms will establish universal, cross-platform characterization protocols.
Acknowledgment
The paper was written as part of the project “Quantum-based Energy Grids (QuGrids)”, which is receiving funding from the programme “Profilbildung 2022”, an initiative of the Ministry of Culture and Science of the State of North Rhine-Westphalia. The contribution of M. R. and D. A. K. has been supported by the Dieter Schwarz Foundation through the Fraunhofer Heilbronn Research and Innovation Centers HNFIZ. Parts of the data used for this study have been generated in the course of propject QORA funded by by Ministerium für Wirtschaft, Arbeit und Tourismus Baden-Württemberg in the frame of the Competence Center Quantum Computing Baden-Württemberg. We acknowledge the use of IBM Quantum services for this work and to advanced services provided by the IBM Quantum Researchers Program. The views expressed are those of the authors, and do not reflect the official policy or position of IBM or the IBM Quantum team.
Appendix A Experimental details
We present the experimental details. Figure 6 shows the optimization convergence across different QAOA circuit sizes on ibmq_kolkata, corresponding to Fig. 5. Table 6 quantifies the structural consistency of our benchmark circuits across different training set sizes. The physical qubits used are identical across different circuit sizes: {12, 15, 18, 21, 23, 24} for the training set and {4, 7, 10, 12, 15, 18, 21, 23, 24} for the prediction set.
Figure 6: Convergence of the optimization process across different circuit sizes on ibmq_kolkata in Fig. 5.
Table 6: Properties of benchmark circuit sets on ibmq_kolkata, analyzed for different numbers of QAOA circuits in Fig. 5.
Training set Prediction set
Property 9 circuits 12 circuits 15 circuits 18 circuits 9 circuits 12 circuits 15 circuits 18 circuits
Depth $27.0\pm 3.5$ $27.0\pm 3.4$ $27.0\pm 3.4$ $27.0\pm 3.4$ $39.0\pm 3.5$ $39.0\pm 3.4$ $39.0\pm 3.4$ $39.0\pm 3.4$
CNOT gates $27.0\pm 10.9$ $27.0\pm 10.7$ $27.0\pm 10.6$ $27.0\pm 10.5$ $78.0\pm 18.6$ $78.0\pm 18.4$ $78.0\pm 18.2$ $78.0\pm 18.1$
Single-qubit gates $50.3\pm 10.8$ $50.3\pm 10.7$ $50.3\pm 10.6$ $50.3\pm 10.5$ $92.3\pm 13.4$ $92.3\pm 13.2$ $92.3\pm 13.1$ $92.3\pm 13.0$
Circuit fidelity $F$ (%) $7.3\pm 8.1 77$$6\pm 8.1 77.$$7\pm 8.1 77.$$7.8\pm 8.1$ $48.3\pm 8.2$ $47.9\pm 8.3$ $47.7\pm 8.3$ $47.5\pm 8.3$
Figure 7(a) compares our parameterized model against the default model using VQE obtained on ibm_kingston in Sec. III.2, while Fig. 7(b) presents results using random circuits on ibmq_kolkata in Sec. III.3.
(a)(b)
Figure 7: Performance comparison of noise models using (a) VQE on ibm_kingston and (b) random circuits on ibmq_kolkata.